# Demystifying Group Relative Policy Optimization: Its Policy Gradient is a U-Statistic

### 03 Mar 2026 v2

@@ 1 Introduction
Since the birth of ChatGPT in 2022, large language models (LLMs) have been increasingly integrated into our work and everyday life. These models are evolving at an extremely face pace and can now perform a broad range of tasks at or beyond human level, ushering in an era of rapid technological advancement across scientific, industrial, and societal domains. One of the underlying technologies that enhance the capabilities of these LLMs is reasoning. The idea is straightforward: because LLMs generate human language, they can be guided to perform multi-step reasoning in a manner similar to humans. Initially, this is often achieved through some simple magical prompts (Wei et al., 2022) – for example, “Let us think step by step” – which encourage the model to decompose complex questions into intermediate reasoning steps and produce more accurate solutions; see Figure 1 for an illustration.
@@
Early progress in LLM reasoning was driven largely by inference-time techniques that avoid retraining the language model, focusing instead on eliciting or searching for more effective intermediate reasoning steps. In parallel, Ouyang et al. (2022) established a scalable reinforcement learning with human feedback (RLHF) pipeline for retraining – more precisely, post-training – LLMs to align their outputs with human values. Its main idea is to (i) collect multiple responses for each prompt and obtain human preference over these responses; (ii) learn a reward model from the resulting preference data (Christiano et al., 2017); and (iii) apply reinforcement learning (RL), specifically the proximal policy optimization algorithm (PPO, Schulman et al., 2017) to fine-tune the LLM parameters so as to maximize the cumulative estimated reward. This work provided a practical demonstration that RL, when combined with human preference feedback, can reliably steer LLM behavior at scale.
@@
This paper addresses these questions by demystifying GRPO from a statistical perspective. Our key observation is that GRPO is deeply connected to U-statistics (Hoeffding, 1948), a connection that has not been explicitly recognized in prior works. Building on this observation, we conduct a comprehensive analysis of GRPO, covering both finite-sample and asymptotic properties, examining both the original algorithm and its subsequent variants, and analyzing both the policy gradient used to update model parameters at each iteration and the suboptimality gap, which measures the difference between the learned policy and the optimal policy; see Figure 3 for a visualization of our theoretical framework.
@@
    To address Q3, we provide finite-sample analyses that characterize the mean squared error (MSE) of the GRPO policy gradient (Theorem 2 & Proposition 3) and derive error bounds on its suboptimality gap (Lemma 6). We further establish parameter consistency and derive the asymptotic distribution of the suboptimality gap without requiring parameter identifiablity (Theorem 7). This result is novel in two respects: (i) existing literature primarily focuses on error bounds for the suboptimality gap, which characterize its order but are not as accurate as its distribution; and (ii) classical asymptotic analyses rely on parameter identifiability, an assumption clearly violated in overparameterized LLMs.
@@
    Finally, to address Q4, we derive a scaling law that delineates how GRPO’s performance depends on the number of sampled outputs per group, and identifies the optimal group size that maximizes its performance (Theorem 8). Notably, this optimal choice depends only on the training data and model architecture, and is independent of other factors such as the training budget or number of iterations. This universality makes our scaling law particularly appealing, as it does require retuning when these factors change. Its universality is further validated empirically (Table 2 & Figure 5).
@@ 2 Related works
Our work connects two modern areas in artificial intelligence – reinforcement learning, and its emerging application to LLM reasoning through RLVR – with the classical theory of U-statistics.
@@ 2.1 Reinforcement learning
    RLHF: Following the seminal work by Ouyang et al. (2022), the field has seen an unprecedented proliferation of literature dedicated to RLHF Christiano et al. (2017), with the goal of aligning the output of LLMs with human preferences through RL (e.g., Munos et al., 2023; Rafailov et al., 2023; Wu et al., 2024, 2025b; Zeng et al., 2024). Theoretically, existing works have explored both the asymptotic distribution of the parameter estimates (Liu et al., 2024) and non-asymptotic error bounds for the sub-optimality gap (e.g., Chowdhury et al., 2024; Zhong et al., 2024; Aminian et al., 2025; Ye et al., 2025; Xu et al., 2025a) and regret (Zhang et al., 2025d). Despite these analyses, GRPO – a key driver of recent breakthroughs in LLM reasoning – remains largely under-theorized.
Finally, RL is closely related to two branches of research in statistics, primarily motivated by healthcare applications; see Chakraborty and Moodie (2013); Kosorok and Laber (2019); Tsiatis et al. (2019); Li et al. (2023); Shi (2025); Ge et al. (2025); Gazi et al. (2026) for reviews. (i) Early works develop RL algorithms to learn optimal dynamic treatment regimes (DTRs), which are individualized strategies that tailor medical interventions to a patient’s unique characteristics and evolving clinical state. Some representative algorithms include Q-learning (Qian and Murphy, 2011; Song et al., 2015), A-learning (Murphy, 2003; Robins, 2004; Shi et al., 2018) and policy-based methods (Zhang et al., 2013; Zhao et al., 2015). These studies focus on short-horizon settings, where treatment decisions are made at a single stage or over a small, finite number of stages. (ii) More recently, the literature has expanded to study MDPs under long-horizon settings (e.g., Ertefaie and Strawderman, 2018; Luckett et al., 2020; Liao et al., 2022; Chen et al., 2024; Li et al., 2024a; Shi et al., 2024a, b; Zhou et al., 2024; Ma et al., 2025; Li et al., 2025b; Zhong et al., 2025) as well as RLHF (e.g., Lee et al., 2024; Liu et al., 2025b, a; Xiao et al., 2025b).
@@ 2.2 Reinforcement learning from verifiable rewards
RLVR, an LLM post-training strategy introduced by (Lambert et al., 2025), directly optimizes LLMs against verifiable outcomes. Unlike RLHF, which requires to learn a reward function from subjective human preferences across multiple candidate responses, RLVR leverages objective feedback signals – for example, by verifying whether a model’s answer matches a ground-truth mathematical solution or by executing the model’s generated code in programming tasks. The original algorithm in Lambert et al. (2025) relied on PPO for policy learning, which learns a separate critic model that evaluates the quality of the learning policy.
GRPO, a major breakthrough in RLVR, drastically scales the reasoning capacities of existing LLMs. Its key ingredient lies in completely eliminating the critic model, sampling multiple reasoning traces, and using their average reward as a proxy for the critic (see Section 3 for details). This approach, originally introduced in DeepSeekMath Shao et al. (2024), was largely popularized by DeepSeek-R1 Guo et al. (2025) and later validated by a series of pioneering open-source reasoning models like Qwen2.5 Yang et al. (2025). Together, these works have sparked a surge of follow-up RLVR algorithms (see, e.g., Zhang et al., 2025a, for a recent review), which can be broadly categorized into three types:
@@
    The second line of works considered different optimization objectives, including entropy-regularized objectives that encourage exploration to prevent the model from entropy collapse (Zhang et al., 2025b; Cheng et al., 2025; Chen et al., 2025), length- or difficulty-aware objectives that balance correctness with concise reasoning (Zhang and Zuo, 2025; Dai et al., 2025), risk-sensitive targets (Ren et al., 2026) and objectives that guide the model to adaptively select reasoning formats (Wu et al., 2025a).
@@
    The third line of works focused on improving GRPO’s training efficiency, either statistically or computationally (Zhang et al., 2025c; Yu et al., 2025; Lin et al., 2025; Xu et al., 2025b). For example, Yan et al. (2025) leveraged off-policy data from stronger models to provide informative reasoning trajectories and enhance learning. Li et al. (2025a) and Zhan et al. (2026) reused previously generated trajectories instead of discarding all old samples after each policy update. Zheng et al. (2025b) explored multiple reasoning paths concurrently to accelerate reasoning speed, while Xu and Ding (2026) generated only a single sampled output per prompt to reduce computational cost.
@@ 3.1 Problem setup
In natural language processing, text is represented as a sequence of discrete units known as tokens. Let $\mathcal{V}$ denote the vocabulary consisting of all such tokens. LLMs produce text via autoregressive next token prediction. Specifically, given an input prompt $X$ (e.g., a user query) represented as a sequence of tokens, the model generates the first token $Y_{1}\in\mathcal{V}$ sampled from a conditional distribution $\pi_{\theta}(\bullet|X)=\mathbb{P}(Y_{1}=\bullet|X)$, parameterized by $\theta\in\Theta$. Next, the second token $Y_{2}$ is produced according to $\pi_{\theta}(\bullet|X,Y_{1})$. At each time step $t$, the model generates $Y_{t}$ according to $\pi_{\theta}(\bullet|X,Y_{<t})$, where $Y_{<t}=(Y_{1},\dots,Y_{t-1})$ represents the previously generated prefix. This iterative procedure continues until the model generates a complete output $Y=(Y_{1},\dots,Y_{T})$, terminating with an “end-of-sequence” (EOS) token $Y_{T}$. To handle conditioning sets of varying lengths, a Transformer architecture (Vaswani et al., 2017) is commonly employed to parameterize $\pi_{\theta}$. For reasoning tasks, the output $Y$ contains both a reasoning trace and a final solution (denoted by $S$); see Figure 1 for an illustration.
@@
Alternatively, this problem can be viewed from a different perspective by treating the entire output sequence $Y$ as a single action. This effectively collapses the time horizon $T$ to $1$. Consequently, the problem is recast as a bandit problem (e.g., Lai and Robbins, 1985), where the data is summarized by the context-action-reward $(X,Y,Z)$ tuples.
@@ 3.2 A meta-algorithm
$g(\theta):=\nabla_{\theta}[\mathbb{E}^{\pi_{\theta}}(Z)]=\mathbb{E}\Big[\sum_{t}\nabla_{\theta}\log_{\pi_{\theta}}(Y_{t}|X,Y_{<t})Z\Big].$ (3)
@@
Equation (3) gives rise to the classical REINFORCE algorithm (Williams, 1992). At the $i$th iteration, the algorithm (i) samples a prompt $X$ from a reference distribution $f(\bullet)$ [^1] While the true distribution $f$ is unknown, it is typically approximated by the empirical distribution of a finite set of prompts. In such cases, our established theoretical guarantees are understood to be conditional on this specific prompt set., (ii) utilizes the current parameter $\theta_{i}$ to generate an output $Y$, (iii) computes its reward $Z$, (iv) constructs a plug-in estimator $\widehat{g}(\theta_{i})=\sum_{t}\nabla_{\theta}\log_{\pi_{\theta}}(Y_{t}|X,Y_{<t})Z$ based on these samples, and (v) updates the parameters via stochastic gradient ascent $\theta_{i+1}\leftarrow\theta_{i}+\eta_{i}\widehat{g}(\theta_{i})$.
While REINFORCE’s gradient estimator is unbiased, it is notoriously susceptible to high variance. A common trick in the literature to reduce its variance is to subtract a baseline term $C_{i}$ (which can vary across iterations $i$) from the reward $Z$. The rationale lies in that the score function $\sum_{t}\nabla_{\theta}\log\pi_{\theta}(Y_{t}|X,Y_{<t})$ has an expectation of zero. Consequently, provided that $C_{i}$ is a function exclusively of the random variable $X$ (being conditionally independent of $Y$ and $Z$ given $X$), the expectation in (3) remains invariant when $Z$ is replaced by its “centered” version, $Z-C_{i}$. As such, the resulting gradient estimator remains unbiased. However, a proper choice of $C_{i}$ can effectively reduce the variance of the estimator.
The most widely adopted $C_{i}$ is the value function $V^{\pi_{\theta_{i}}}(X)=\mathbb{E}^{\pi_{\theta_{i}}}(Z|X)$. This choice gives rise to advantage actor-critic (A2C) algorithms (e.g., Mnih et al., 2016), which, in addition to the policy (the actor), learn a value function (the critic) to estimate the baseline and calculate the advantage function (the difference between the return and the critic) for variance reduction. As mentioned in the introduction, this critic network is also a core component of PPO. However, in reasoning tasks, maintaining and updating a separate critic network is extremely computationally intensive. GRPO addresses this inefficiency by eliminating the critic network entirely. It samples multiple outputs for each prompt and utilizes their group mean as the baseline term (see Equation (4)). Because sampling from the policy is drastically more efficient than training a secondary critic model, GRPO offers a highly scalable solution for LLM reasoning. In Section 4, we show that this solution is also mathematically elegant.
@@
We conclude this section by noting that while we focus on analyzing Algorithm 1 for building intuition, it departs from standard production-level implementations in three aspects: (i) it omits the reward normalization commonly used in GRPO; (ii) it does not incorporate the importance sampling strategy used in GRPO to enhance sample efficiency through gradient updates per batch; (iii) it excludes the Kullback–Leibler (KL) divergence penalty, which prevents the policy from deviating excessively from the reference model. These simplifications were implemented in GRPO variants such as Dr-GRPO (Liu et al., 2025c) and GPG (Chu et al., 2025), which we analyze rigorously in Sections 4.1 and 4.2. In Section LABEL:sec:ext of the Supplementary Material, we will further bridge these gaps to align our theoretical analysis with standard GRPO implementations.

### 22 Mar 2026 v3

@@ 1 Introduction
Since the birth of ChatGPT in 2022, large language models (LLMs) have been increasingly integrated into our work and everyday life. These models are evolving at an extremely fast pace and can now perform a broad range of tasks at or beyond human level, ushering in an era of rapid technological advancement across scientific, industrial, and societal domains. One of the underlying technologies that enhance the capabilities of these LLMs is reasoning. The idea is straightforward: because LLMs generate human language, they can be guided to perform multi-step reasoning in a manner similar to humans. Initially, this is often achieved through some simple magical prompts (Wei et al., 2022) – for example, “Let us think step by step” – which encourage the model to decompose complex questions into intermediate reasoning steps and produce more accurate solutions; see Figure 1 for an illustration.
@@
Early progress in LLM reasoning was driven largely by inference-time techniques that avoid retraining the language model, focusing on eliciting or searching for more effective intermediate reasoning steps. In parallel, Ouyang et al. (2022) established a scalable reinforcement learning with human feedback (RLHF) pipeline for retraining – more precisely, post-training – LLMs to align their outputs with human values. Its main idea is to (i) collect multiple responses for each prompt and obtain human preference over these responses; (ii) learn a reward model from the resulting preference data (Christiano et al., 2017); and (iii) apply reinforcement learning (RL), specifically the proximal policy optimization algorithm (PPO, Schulman et al., 2017) to fine-tune the LLM parameters so as to maximize the cumulative estimated reward. This work provided a practical demonstration that RL, when combined with human preference feedback, can reliably steer LLM behavior at scale.
@@
This paper addresses these questions by demystifying GRPO from a statistical perspective. Our key observation is that GRPO is deeply connected to U-statistics (Hoeffding, 1948), a connection that has not been explicitly recognized in prior works. Building on this observation, we conduct a comprehensive analysis of GRPO, covering both finite-sample and asymptotic properties, examining both the original algorithm and its variants, and analyzing both the policy gradient estimator used to update model parameters at each iteration and the suboptimality gap, which measures the difference between the learned policy and the optimal policy; see Figure 3 for a visualization of our theoretical framework.
@@
    To address Q3, we provide finite-sample analyses that characterize the mean squared error (MSE) of the GRPO policy gradient (Theorem 2 & Proposition 3) and derive error bounds on its suboptimality gap (Lemma 6). We further establish parameter consistency and derive the asymptotic distribution of the suboptimality gap without requiring parameter identifiability (Theorem 8). This result is novel in two respects: (i) existing literature primarily focuses on error bounds for the suboptimality gap, which characterize its order but are not as accurate as its distribution; and (ii) classical asymptotic analyses rely on parameter identifiability, an assumption clearly violated in overparameterized LLMs.
@@
    Finally, to address Q4, we derive a scaling law that delineates how GRPO’s performance depends on the number of sampled outputs per group, and identifies the optimal group size that maximizes its performance (Theorem 7). Notably, this optimal choice depends only on the training data and model architecture, and is independent of other factors such as the training budget or number of iterations. This universality makes our scaling law particularly appealing, as it does not require retuning when these factors change. Its universality is further validated empirically (Table 2 & Figure 5).
@@ 2 Related works
Our work connects two modern areas in artificial intelligence – reinforcement learning (Section 2.1), and its emerging application to LLM reasoning through RLVR (Section 2.2) – with the classical theory of U-statistics (Section 2.3).
@@ 2.1 Reinforcement learning
    RLHF: Following the seminal work by Ouyang et al. (2022), the field has seen an unprecedented proliferation of literature dedicated to RLHF (Christiano et al., 2017), with the goal of aligning the output of LLMs with human preferences through RL (e.g., Munos et al., 2023; Rafailov et al., 2023; Wu et al., 2024, 2025b; Zeng et al., 2024). Theoretically, existing works have explored both the asymptotic distribution of the parameter estimates (Liu et al., 2024) and finite-sample error bounds for the sub-optimality gap (e.g., Chowdhury et al., 2024; Zhong et al., 2024; Aminian et al., 2025; Ye et al., 2025; Xu et al., 2025a) and regret (Zhang et al., 2025d). Despite these analyses, GRPO – a key driver of recent breakthroughs in LLM reasoning – remains largely under-theorized.
Finally, RL is closely related to two branches of research in statistics, primarily motivated by healthcare applications; see Chakraborty and Moodie (2013); Kosorok and Laber (2019); Tsiatis et al. (2019); Li et al. (2023); Shi (2025); Ge et al. (2025); Gazi et al. (2026) for reviews. (i) Early works develop RL algorithms to learn optimal dynamic treatment regimes (DTRs), which are individualized strategies that tailor medical interventions to a patient’s unique characteristics and evolving clinical state. Some representative algorithms include Q-learning (Qian and Murphy, 2011; Song et al., 2015), A-learning (Murphy, 2003; Robins, 2004; Shi et al., 2018) and policy-based methods (Zhang et al., 2013; Zhao et al., 2015). These studies focus on short-horizon settings, where treatment decisions are made at a single stage or over a small, finite number of stages. (ii) More recently, the literature has expanded to study MDPs under long-horizon settings (e.g., Ertefaie and Strawderman, 2018; Luckett et al., 2020; Liao et al., 2022; Chen et al., 2024; Li et al., 2024a; Shi et al., 2024a, b; Zhou et al., 2024; Ma et al., 2025; Jin et al., 2025; Li et al., 2025b; Zhong et al., 2025) as well as RLHF (e.g., Lee et al., 2024; Liu et al., 2025b, a; Lu et al., 2025; Xiao et al., 2025b).
@@ 2.2 Reinforcement learning from verifiable rewards
RLVR, an LLM post-training strategy introduced by Lambert et al. (2025), directly optimizes LLMs against verifiable outcomes. Unlike RLHF, which requires to learn a reward function from subjective human preferences across multiple candidate responses, RLVR leverages objective feedback signals – for example, by verifying whether a model’s answer matches a ground-truth mathematical solution or by executing the model’s generated code in programming tasks. The original algorithm in Lambert et al. (2025) relied on PPO for policy learning, which learns a separate critic model that evaluates the quality of the learning policy.
GRPO, a major breakthrough in RLVR, drastically scales the reasoning capacities of existing LLMs. Its key ingredient lies in completely eliminating the critic model, sampling multiple reasoning traces, and using their average reward as a proxy for the critic (see Section 3 for details). This approach, originally introduced in DeepSeekMath (Shao et al., 2024), was largely popularized by DeepSeek-R1 (Guo et al., 2025) and later validated by a series of pioneering open-source reasoning models like Qwen2.5 (Yang et al., 2025). Together, these works have sparked a surge of follow-up RLVR algorithms (see, e.g., Zhang et al., 2025a, for a recent review), which can be broadly categorized into three types:
@@
    The second line of works considered different optimization objectives, including entropy-regularized objectives that encourage exploration to prevent the model from entropy collapse (Zhang et al., 2025b; Cheng et al., 2025; Chen et al., 2025), length- or difficulty-aware objectives that balance correctness with concise reasoning (Zhang and Zuo, 2025; Dai et al., 2025a), risk-sensitive targets (Ren et al., 2026) and objectives that guide the model to adaptively select reasoning formats (Wu et al., 2025a).
@@
    The third line of works focused on improving GRPO’s training efficiency, either statistically or computationally (Dai et al., 2025b; Yu et al., 2025; Lin et al., 2025; Xu et al., 2025b; Zhang et al., 2025c). For example, Yan et al. (2025) leveraged off-policy data from stronger models to provide informative reasoning trajectories and enhance learning. Li et al. (2025a) and Zhan et al. (2026) reused previously generated trajectories instead of discarding all old samples after each policy update. Zheng et al. (2025b) explored multiple reasoning paths concurrently to accelerate reasoning speed, while Xu and Ding (2026) generated only a single sampled output per prompt to reduce computational cost.
@@ 3.1 Problem setup
In natural language processing, text is represented as a sequence of discrete units known as tokens. Let $\mathcal{V}$ denote the vocabulary consisting of all such tokens. LLMs produce text via autoregressive next token prediction. Specifically, given an input prompt $X$ (e.g., a user query) represented as a sequence of tokens, the model generates the first token $Y_{1}\in\mathcal{V}$ sampled from a conditional distribution $\pi_{\theta}(\bullet|X)=\mathbb{P}(Y_{1}=\bullet|X)$, parameterized by $\theta\in\Theta$. Next, the second token $Y_{2}$ is produced according to $\pi_{\theta}(\bullet|X,Y_{1})$. At each time step $t$, the model generates $Y_{t}$ according to $\pi_{\theta}(\bullet|X,Y_{<t})$, where $Y_{<t}=(Y_{1},\dots,Y_{t-1})$ represents the previously generated prefix. This iterative procedure continues until the model generates a complete output $Y=(Y_{1},\dots,Y_{T})$, terminating with an “end-of-sequence” token $Y_{T}$. To handle conditioning sets of varying lengths, a Transformer architecture (Vaswani et al., 2017) is commonly employed to parameterize $\pi_{\theta}$. For reasoning tasks, the output $Y$ contains both a reasoning trace and a final solution (denoted by $S$); see Figure 1 for an illustration.
@@
Alternatively, this problem can be viewed from a different perspective by treating the entire output sequence $Y$ as a single action. This effectively collapses the time horizon $T$ to 11. Consequently, the problem is recast as a bandit problem (e.g., Lai and Robbins, 1985), where the data is summarized by the context-action-reward $(X,Y,Z)$ tuples.
@@ 3.2 A meta-algorithm
$g(\theta):=\nabla_{\theta}[\mathbb{E}^{\pi_{\theta}}(Z)]=\mathbb{E}\Big[\sum_{t}\nabla_{\theta}\log\pi_{\theta}(Y_{t}|X,Y_{<t})Z\Big].$ (3)
Equation (3) gives rise to the classical REINFORCE algorithm (Williams, 1992). At the $i$th iteration, the algorithm (i) samples a prompt $X$ from a reference distribution $f(\bullet)$ [^1] While the true distribution $f$ is unknown, it is typically approximated by the empirical distribution of a finite set of prompts. In such cases, our established theoretical guarantees are understood to be conditional on this specific prompt set., (ii) utilizes the current parameter $\theta_{i}$ to generate an output $Y$, (iii) computes its reward $Z$, (iv) constructs a plug-in estimator $\widehat{g}(\theta_{i})=\sum_{t}\nabla_{\theta}\log\pi_{\theta}(Y_{t}|X,Y_{<t})Z$ based on these samples, and (v) updates the parameters via stochastic gradient ascent $\theta_{i+1}\leftarrow\theta_{i}+\eta_{i}\widehat{g}(\theta_{i})$ with learning rate $\eta_{i}>0$.
@@
While REINFORCE’s gradient estimator is unbiased, it is notoriously susceptible to high variance. A common trick in the literature to reduce its variance is to subtract a baseline term $C_{i}$ (which can vary across iterations ii) from the reward $Z$. The rationale lies in that the score function $\sum_{t}\nabla_{\theta}\log\pi_{\theta}(Y_{t}|X,Y_{<t})$ has an expectation of zero. Consequently, provided that $C_{i}$ is a function exclusively of the random variable $X$ (being conditionally independent of $Y$ and $Z$ given $X$), the expectation in (3) remains invariant when $Z$ is replaced by its “centered” version, $Z-C_{i}$. As such, the resulting gradient estimator remains unbiased. However, a proper choice of $C_{i}$ can effectively reduce the variance of the estimator.
The most widely adopted $C_{i}$ is the value function $V^{\pi_{\theta_{i}}}(X)=\mathbb{E}^{\pi_{\theta_{i}}}(Z|X)$. This choice gives rise to advantage actor-critic (A2C) algorithms (e.g., Mnih et al., 2016), which, in addition to the policy (the actor), learn a value function (the critic) to estimate the baseline and calculate the advantage function (the difference between the return and the critic) for variance reduction. However, in reasoning tasks, maintaining and updating a separate critic network is extremely computationally intensive. GRPO addresses this inefficiency by eliminating the critic network entirely. It samples multiple outputs for each prompt and utilizes their group mean as the baseline term (see Equation (4)). Because sampling from the policy is drastically more efficient than training a critic model, GRPO offers a highly scalable solution to LLM reasoning. In Section 4, we show that this solution is also mathematically elegant.
@@
We conclude this section by noting that while we focus on analyzing Algorithm 1 for building intuition, it simplifies the standard production-level implementation in three aspects: (i) it omits the reward normalization used in GRPO; (ii) it does not incorporate the importance sampling strategy used in GRPO to enhance sample efficiency through multiple gradient updates per batch; (iii) it excludes the Kullback–Leibler (KL) divergence penalty, which prevents the policy from deviating excessively from the reference model. These simplifications were implemented in GRPO variants such as Dr-GRPO (Liu et al., 2025c) and GPG (Chu et al., 2025), which we analyze rigorously in Sections 4.1 and 4.2. In Section A of the Supplementary Material, we will bridge these gaps to align our theoretical analyses with standard GRPO implementations.